# KnowTuning: Knowledge-aware Fine-tuning for Large Language Models

### v2

@@ Abstract
Yougang Lyu1,3   Lingyong Yan2   Shuaiqiang Wang2  Haibo Shi2  Dawei Yin2 Pengjie Ren1  Zhumin Chen1  Maarten de Rijke3   Zhaochun Ren4thanks:  Corresponding author. 1Shandong University, Qingdao, China   2Baidu Inc., Beijing, China 3University of Amsterdam, Amsterdam, The Netherlands 4Leiden University, Leiden, The Netherlands youganglyu@gmail.com, {yanlingyong, wangshuaiqiang, shihaibo}@baidu.com yindawei@acm.org, jay.ren@outlook.com, chenzhumin@sdu.edu.cn m.derijke@uva.nl, z.ren@liacs.leidenuniv.nl
@@ 1 Introduction
  • We focus on improving the fine-grained and coarse-grained knowledge awareness of LLMs via fine-tuning for knowledge-intensive tasks.
@@
  • We demonstrate the effectiveness of KnowTuning in the generic and medical domain QA datasets through automatic and human evaluations, across various sizes of LLMs. Furthermore, KnowTuning generates more facts with less factual error rate under fine-grained facts evaluation.
@@ 2.1 LLMs for knowledge-intensive Tasks
LLM have been applied to various knowledge-intensive tasks Moiseev et al. (2022); Yu et al. (2023b); Khattab et al. (2022); Tian et al. (2023); Zhang et al. (2023a); Xu et al. (2023b); Mishra et al. (2023); Nguyen et al. (2023). Previous work mainly focus on knowledge-intensive tasks with short-form answers.  Liu et al. (2022b) use few-shot demonstrations to elicit relevant knowledge statements from LLMs for QA tasks. Liu et al. (2022a) train a neural model to generate relevant knowledge through reinforcement learning for QA tasks. Liu et al. (2023a) propose a unified model for generating relevant knowledge and solving QA tasks.
However, these methods primarily address multiple-choice QA, rather than the more complex open-ended knowledge-intensive QA tasks Krishna et al. (2021); Kadavath et al. (2022); Liu et al. (2022a, 2023a); Kang et al. (2023), which aim to solve questions that require detailed explanations and extensive domain knowledge. Recent research indicates that LLMs face challenges in tackling complex knowledge-intensive QA tasks Yu et al. (2023a); Bai et al. (2023); Chang et al. (2023). In particular, they are prone to generating responses that are non-factual Lee et al. (2022); Sun et al. (2023); Su et al. (2022), incomplete Singhal et al. (2022); Bian et al. (2024), or illogical Chen et al. (2023b); Zhong et al. (2023). Recently, for open-ended knowledge-intensive tasks,  Tian et al. (2023) propose a method FacTune to improve factuality. Specifically, they first automatically evaluate the proportion of correct facts in candidate answers as factuality scores, and fine-tuning LLMs to increase the likelihood of generating answers with higher factuality scores. In contrast, our work focus on improving the knowledge awareness of LLMs at multiple essential aspects simultaneously, for solving complex knowledge-intensive QA tasks.
@@ 2.2 Fine-tuning for LLMs
Another line of fine-tuning methods fuse information about the quality of the generated answers into the supervision signals Zhao et al. (2023); Guo et al. (2023); Wang et al. (2023d); Dong et al. (2023); Chen et al. (2024).  Rafailov et al. (2023) propose direct preference optimization (DPO) to directly optimize LLMs on the pair-wise comparison set. Song et al. (2023) propose Preference Ranking Optimizatio (PRO) to fine-tune LLMs on list-wise comparison sets.  Yuan et al. (2023) propose a margin-rank loss to optimize the LLMs on comparison sets. Since collecting large-scale human judgment for the quality of generated answers is expensive, Bai et al. (2022) and  Lee et al. (2023) propose reinforcement learning from AI feedback (RLAIF) methods to leverage off-the-shelf LLMs to annotate general helpfulness scores. In contrast, our work focuses on enhancing the fine-grained and coarse-grained knowledge-awareness of LLMs to improve performance in terms of completeness, factuality, and logicality simultaneously.
@@ 3.3 Coarse-grained Knowledge Comparison
Finally, we combine the rephrased answer $a^{r}_{i}$ and the illogical answer $a^{l}_{i}$ into knowledge logicality comparison set as follows:
@@ 3.4 Training
To improve the knowledge awareness of LLMs for solving complex knowledge-intensive tasks, KnowTuning includes fine-grained knowledge augmentation training and coarse-grained knowledge comparison training. Specifically, we first train LLMs on fine-grained knowledge augmentation dataset $\mathcal{D}_{ka}$, resulting in a model denoted as $\pi_{ka}$. Then, KnowTuning aims to further improve the coarse-grained knowledge awareness of the model $\pi_{ka}$ in completeness, factuality, and logicality. To accomplish this, we rewrite the DPO Rafailov et al. (2023) loss to obtain the coarse-grained knowledge comparison loss as follows:
$\begin{split}\mbox{}\mathcal{L}_{kc}\!={}&\!\mathbb{E}_{(q,(a_{w},a_{l}))\sim\mathcal{D}_{kc}}\bigg{[}\!\log\sigma\bigg{(}\!\beta\log\!\frac{\pi_{kc}(a_{w}|q)}{\pi_{ka}(a_{w}|q)}\mbox{}\\ &-\beta\log\frac{\pi_{kc}(a_{l}|q)}{\pi_{ka}(a_{l}|q)}\bigg{)}\bigg{]},\end{split}$ (18)
where $(a_{w},a_{l})$ denotes the answer pair of the question $q\in D_{kc}$, and $a_{w}$ is the better answer.
@@ 4.1 Research questions
4.2 Datasets
More details of datasets are in Appendix A.
4.3 Baselines
We compare our model with the following baselines:
More details of baselines are in Appendix B.
@@ 4.4 Evaluation Metrics
We present our experimental results using two evaluation metrics: automatic evaluation and human-based evaluation. Since ROUGE ROUGE (2004) and BLEU Papineni et al. (2002) can not effectively evaluate the quality of answers for complex questions Krishna et al. (2021); Xu et al. (2023a), recent studies propose to use GPT-4 for evaluating the quality of LLMs answers Zheng et al. (2024); Dubois et al. (2023); Fu et al. (2023). Consequently, given golden label as a reference, we employ GPT-4 to rate generated answers on three aspects: completeness, factuality, and logicality, on a range of 1 to 10. Following Singhal et al. (2022); Zheng et al. (2024); Zhang et al. (2023b), we define completeness, factuality and logicality as: (i) Completeness: it examines whether the answers provide comprehensive and sufficient knowledge to the questions. (ii) Factuality: it examines whether the knowledge in the answers is factually correct. (iii) Logicality: it examines whether the knowledge in the answers is logically structured. Following Li et al. (2023); Chen et al. (2023a), we define “Win-Tie-Lose” as: (i) Win: KnowTuning wins twice, or wins once and ties once. (ii) Tie: KnowTuning ties twice, or wins once and loses once. (iii) Lose: KnowTuning loses twice, or loses once and ties once.
@@ 4.5 Implementation details
KnowTuning vs FactTune Dolly 58.00 11.00 31.00 32.50 66.50 11.00 35.00 53.00 12.00 +23.83
@@ 5.1 Main results (RQ1)
Automatic evaluation. Table 1 presents the reference-based GPT-4 evaluation results for both generic and medical domain QA datasets. Across all metrics, KnowTuning outperforms the baseline models in these domains. Based on the results, we have two main observations:
  • KnowTuning consistently surpasses baselines in terms of completeness, factuality and logicality, across generic and domain-specific QA datasets. Compared with Base and SFT, KnowTuning focuses on improving fine-grained and coarse-grained knowledge awareness of LLMs, which significantly improves the performance. Compared with RLAIF and FactTune, KnowTuning is more effective in improving the performance of LLMs on complex knowledge-intensive QA in multiple aspects. The reason is that RLAIF improves the performance by calculating overall helpfulness scores and FactTune focuses on improving the factuality, they ignore improving the knowledge awareness of LLMs in multiple essential aspects simultaneously.
@@
Table 4: Ablation study evaluated by GPT-4 on the generic QA dataset. The backbone model is Llama2-7b-base.
5.2 Fine-grained facts evaluation (RQ2)
@@ 6 Conclusions
In this paper, we focus on improving the knowledge awareness of LLMs via fine-tuning for complex knowledge-intensive tasks. We have proposed KnowTuning to fine-tune LLMs through fine-grained knowledge augmentation and coarse-grained knowledge comparison stages. We have conducted comprehensive experiments on generic and medical domain QA datasets, demonstrating the effectiveness of KnowTuning through automatic and human evaluations, across various sizes of LLMs. Moreover, KnowTuning generates more facts with less factual error rate under fine-grained facts evaluation. Our code and dataset are available at https://anonymous.4open.science/r/KnowTuning-345D.
@@
In this study, KnowTuning is mainly aimed at generic and medical knowledge-intensive tasks, we plan to adopt KnowTuning to other tasks such as legal domain QA Zhong et al. (2020) and mathematical reasoning Luo et al. (2023). Moreover, our efforts have been concentrated on enhancing the knowledge awareness of LLMs during the fine-tuning stage. Future studies will aim to explore improving knowledge awareness of LLMs in the pre-training stage Rosset et al. (2020).
Ethics Statement
KnowTuning mainly focuses on completeness, factuality, and logicality, but not social bias or the potential for generating harmful or toxic content Hewitt et al. (2024). We plan to adopt our method to reduce social bias and harmful content at fine-grained and coarse-grained levels in future work.
@@ Appendix B Details of Baselines
  • FactTune Tian et al. (2023): We follow Min et al. (2023) to first break eacn candidate answers into individual facts, and prompt LLMs to measure the correctness of each fact based on the golden answer as a reference. [^1] https://github.com/shmsw25/FActScore Then, we construct factuality comparison sets by the percentage of correct facts. Finally, we adpot DPO Rafailov et al. (2023) for factuality comparison sets optimization.
@@ C.1 GPT-4 Evaluation
This section provides specifics of the GPT-4 prompt utilized for reference-based evaluation, employing gpt4-turbo. Figure 3 illustrates the adapted prompt from  Zheng et al. (2024), aimed at assessing the completeness, factuality, and logicality of answers. To avoid positional bias Ko et al. (2020); Wang et al. (2023e), we evaluate each answer in both positions during two separate runs.
@@ C.3 Fine-grained facts evaluation
Following Min et al. (2023), we first break candidate answers into individual facts, and use gpt-3.5-turbo to measure the correctness of each fact based on the golden answer as a reference.1
@@ D.3 Training
During the training phase, the AdamW optimizer Loshchilov and Hutter (2019) is utilized with initial learning rates of $1\cdot 10^{-4}$ for SFT and $5\cdot 10^{-6}$ for DPO. The batch sizes for SFT and DPO are set to 32 and 16, respectively, with SFT undergoing 3 epochs of training and DPO 1 epoch. The filtering and deleting percentages, $\alpha$ and $\beta$, are both fixed at 0.5. We determine the hyperparameters through pilot experiments. Training leverages PEFT Mangrulkar et al. (2022), LLaMA-Factory Hiyouga (2023) and LoRA Hu et al. (2022).

### v3

@@ Abstract
Yougang Lyu1,3   Lingyong Yan2   Shuaiqiang Wang2  Haibo Shi2  Dawei Yin2 Pengjie Ren1  Zhumin Chen1  Maarten de Rijke3   Zhaochun Ren4thanks:  Corresponding author. 1Shandong University, Qingdao, China   2Baidu Inc., Beijing, China 3University of Amsterdam, Amsterdam, The Netherlands 4Leiden University, Leiden, The Netherlands youganglyu@gmail.com, {yanlingyong, wangshuaiqiang}@baidu.com haiboshi@outlook.com, yindawei@acm.org, jay.ren@outlook.com chenzhumin@sdu.edu.cn, m.derijke@uva.nl, z.ren@liacs.leidenuniv.nl
@@ 1 Introduction
  • We focus on systematically enhancing the knowledge awareness of LLMs at both fine-grained and coarse-grained levels to address complex knowledge-intensive tasks.
@@
  • We demonstrate the effectiveness of KnowTuning in the generic and medical domain QA datasets through automatic and human evaluations, across various sizes of LLMs. Furthermore, KnowTuning generates more facts with less factual error rate under fine-grained facts evaluation. [^1] The code is available at https://github.com/youganglyu/KnowTuning
@@ 2.1 LLMs for Knowledge-intensive Tasks
LLM have been applied to various knowledge-intensive tasks Moiseev et al. (2022); Yu et al. (2023b); Khattab et al. (2022); Tian et al. (2023); Zhang et al. (2023a); Xu et al. (2023b); Mishra et al. (2023); Nguyen et al. (2023); Zhang et al. (2024). Previous work mainly focus on knowledge-intensive tasks with short-form answers.  Liu et al. (2022b) use few-shot demonstrations to elicit relevant knowledge statements from LLMs for QA tasks. Liu et al. (2022a) train a neural model to generate relevant knowledge through reinforcement learning for QA tasks. Liu et al. (2023a) propose a unified model for generating relevant knowledge and solving QA tasks.
However, these methods primarily address multiple-choice QA, rather than the more complex open-ended knowledge-intensive QA tasks Krishna et al. (2021); Kadavath et al. (2022); Liu et al. (2022a, 2023a); Kang et al. (2023), which aim to solve questions that require detailed explanations and extensive domain knowledge. Recent research indicates that LLMs face challenges in tackling complex knowledge-intensive QA tasks Yu et al. (2023a); Bai et al. (2023); Chang et al. (2023). In particular, they are prone to generating responses that are non-factual Lee et al. (2022); Sun et al. (2023); Su et al. (2022), incomplete Singhal et al. (2022); Bian et al. (2024), or illogical Chen et al. (2023b); Zhong et al. (2023). Recently, for open-ended knowledge-intensive tasks,  Tian et al. (2023) propose a method FacTune to improve factuality. Specifically, they first automatically evaluate the proportion of correct facts in candidate answers as factuality scores, and fine-tuning LLMs to increase the likelihood of generating answers with higher factuality scores. In contrast, we focus on improving the knowledge awareness of LLMs at multiple essential aspects simultaneously, for solving complex knowledge-intensive QA tasks.
@@ 2.2 Fine-tuning for LLMs
Another line of fine-tuning methods fuse information about the quality of the generated answers into the supervision signals Zhao et al. (2023); Guo et al. (2023); Wang et al. (2023d); Dong et al. (2023); Chen et al. (2024); Zhao et al. (2024).  Rafailov et al. (2023) propose direct preference optimization (DPO) to directly optimize LLMs on the pair-wise comparison set. Song et al. (2023) propose preference ranking optimization (PRO) to fine-tune LLMs on list-wise comparison sets.  Yuan et al. (2023) propose a margin-rank loss to optimize the LLMs on comparison sets. Since collecting large-scale human judgment for the quality of generated answers is expensive, Bai et al. (2022) and  Lee et al. (2023) propose reinforcement learning from AI feedback (RLAIF) methods to leverage off-the-shelf LLMs to annotate general helpfulness scores. In contrast, our work focuses on enhancing the fine-grained and coarse-grained knowledge-awareness of LLMs to improve performance in terms of completeness, factuality, and logicality simultaneously.
@@ 3.3 Coarse-grained Knowledge Comparison
Next, we combine the rephrased answer $a^{r}_{i}$ and the illogical answer $a^{l}_{i}$ into knowledge logicality comparison set as follows:
@@ 3.4 Training
To improve the knowledge awareness of LLMs for solving complex knowledge-intensive tasks, KnowTuning includes fine-grained knowledge augmentation training and coarse-grained knowledge comparison training. Specifically, we first train LLMs on fine-grained knowledge augmentation dataset $\mathcal{D}_{ka}$, resulting in a model denoted as $\pi_{ka}$. To improve the coarse-grained knowledge awareness of the model $\pi_{ka}$, we rewrite the DPO Rafailov et al. (2023) loss as follows:
$\mathcal{L}_{dpo}\!={}\!-\mathbb{E}_{(q,(a_{w},a_{l}))\sim\mathcal{D}_{kc}}\bigg{[}\!\log\sigma\bigg{(}\!\beta\log\!\frac{\pi_{kc}(a_{w}|q)}{\pi_{ka}(a_{w}|q)}\mbox{}$
$-\beta\log\frac{\pi_{kc}(a_{l}|q)}{\pi_{ka}(a_{l}|q)}\bigg{)}\bigg{]},$ (18)
where $(a_{w},a_{l})$ denotes the answer pair of the question $q\in D_{kc}$, and $a_{w}$ is the better answer. To maintain coarse-grained knowledge awareness of better answers, we add SFT loss into the coarse-grained knowledge comparison loss:
$\mathcal{L}_{kc}\!=\mathcal{L}_{dpo}+\gamma\mathcal{L}_{\mathrm{sft}},$ (19)
where $\mathcal{L}_{\mathrm{sft}}$ is a term for better answers $a_{w}$ and $\gamma$ is a scalar weighting hyperparameter.
@@ 4.1 Research Questions
Dolly MedQuAD NQ ELI5
Method METEOR BERTScore METEOR BERTScore METEOR BERTScore METEOR BERTScore
Backbone Language Model: Llama2-7b-base
Base 12.29 78.07 12.79 78.44 05.10 72.70 09.09 76.05
SFT 14.01 84.38 19.95 80.97 07.55 76.71 11.96 79.65
RLAIF 17.60 85.31 20.60 83.82 10.77 79.62 13.66 80.41
FactTune 16.84 85.16 21.82 82.99 10.08 79.09 14.19 80.83
KnowTuning 19.56 86.37 24.71 84.28 12.22 80.54 16.32 81.74
Backbone Language Model: Llama2-13b-base
Base 11.59 77.90 12.12 78.29 05.51 73.80 07.79 75.63
SFT 15.31 84.39 19.66 82.34 08.70 78.18 12.00 81.21
RLAIF 19.03 85.43 20.37 83.13 11.79 80.30 13.61 82.06
FactTune 18.59 85.38 21.42 83.49 11.37 80.02 13.74 82.16
KnowTuning 20.01 86.32 25.21 84.41 12.56 80.74 14.45 83.06
Table 1: Lexicon-based and semantic-based evaluation on generic and medical QA. The best performance is highlighted in bold.
4.2 Datasets
To evaluate the performance across a wider range of knowledge-intensive tasks, we further evaluate generic QA models on two representative test sets from knowledge intensive language tasks (KILT) benchmark Petroni et al. (2021):
  • NQ Kwiatkowski et al. (2019) consists of real questions directed to the Google search engine. Every question is paired with a corresponding Wikipedia page that includes a detailed long-form answer and a concise short answer. We filter questions and corresponding long answers as testing QA pairs.
  • ELI5 Fan et al. (2019) includes a set of question-answer-evidence triples. The questions are complex, and the responses are comprehensive, explanatory, and presented in a free-form style. We filter questions and corresponding answers as testing QA pairs.
More details of datasets are in Appendix A.
4.3 Baselines
We compare our model with the following baselines:
More details of baselines are in Appendix B.
@@ 4.4 Evaluation Metrics
We present our experimental results using two evaluation metrics: automatic evaluation and human-based evaluation. Following previous studies Clinciu et al. (2021); Slobodkin et al. (2023), we employ two automatic metrics for absolute quality evaluation: the lexicon-based metric METEOR Banerjee and Lavie (2005) and the semantic-based metric BERTScore Zhang et al. (2019). Since recent studies propose that GPT-4 can effectively evaluate the quality of LLMs answers Zheng et al. (2024a); Dubois et al. (2023); Fu et al. (2023), we also conduct GPT-4 pairwise evaluation. Specifically, given the golden label as a reference, we employ GPT-4 to rate generated answers on three aspects: completeness, factuality, and logicality, on a range of 1 to 10. Following Singhal et al. (2022); Zheng et al. (2024a); Zhang et al. (2023b), we define completeness, factuality and logicality as: (i) Completeness: it examines whether the answers provide comprehensive and sufficient knowledge to the questions. (ii) Factuality: it examines whether the knowledge in the answers is factually correct. (iii) Logicality: it examines whether the knowledge in the answers is logically structured. Following Li et al. (2023); Chen et al. (2023a), we define “Win-Tie-Lose” as: (i) Win: KnowTuning wins twice, or wins once and ties once. (ii) Tie: KnowTuning ties twice, or wins once and loses once. (iii) Lose: KnowTuning loses twice, or loses once and ties once.
@@ 5 Experimental Results and Analysis
KnowTuning vs FactTune Dolly 58.00 11.00 31.00 32.50 56.50 11.00 35.00 53.00 12.00 +23.83
@@ 5.1 Main Results (RQ1)
Automatic evaluation. Table 1 and Table 2 present the reference-based GPT-4 evaluation results and absolute quality evaluation results for both generic and medical domain QA datasets. Across all metrics, KnowTuning outperforms the baseline models in these domains. Based on the results, we have three main observations:
  • KnowTuning demonstrates effectiveness under lexicon-based and semantic-based evaluations. As shown in Table 1, our method consistently improves the absolute quality of answers for general and medical QA tasks. Furthermore, these results illustrate the ability of our method to generalize to a wider range of knowledge-intensive datasets, such as NQ and ELI5.
  • KnowTuning consistently outperforms baselines in terms of completeness, factuality and logicality, across generic and domain-specific QA datasets. Compared with Base and SFT, KnowTuning focuses on improving fine-grained and coarse-grained knowledge awareness of LLMs, which significantly improves the performance. Compared with RLAIF and FactTune, KnowTuning is more effective in improving the performance of LLMs on complex knowledge-intensive QA in multiple aspects. The reason is that RLAIF improves the performance by calculating overall helpfulness scores and FactTune focuses on improving the factuality, they ignore improving the knowledge awareness of LLMs in multiple essential aspects simultaneously.
@@
Table 5: Ablation study evaluated by GPT-4 on the generic QA dataset. The backbone model is Llama2-7b-base. -KA indicates the exclusion of fine-grained knowledge augmentation, -KCC indicates the exclusion of completeness comparison, -KFC indicates the exclusion of factuality comparison, -KLC indicates the exclusion of logicality comparison, and -KC indicates the exclusion of all coarse-grained knowledge comparisons.
5.2 Fine-grained Fact Evaluation (RQ2)
@@ 6 Conclusions
In this paper, we focus on improving the knowledge awareness of LLMs via fine-tuning for complex knowledge-intensive tasks. We have proposed KnowTuning to fine-tune LLMs through fine-grained knowledge augmentation and coarse-grained knowledge comparison stages. We have conducted comprehensive experiments on generic and medical domain QA datasets, demonstrating the effectiveness of KnowTuning through automatic and human evaluations, across various sizes of LLMs. Moreover, KnowTuning generates more facts with less factual error rate under fine-grained facts evaluation.
@@
In this study, KnowTuning is mainly aimed at generic and medical knowledge-intensive tasks, we plan to adopt KnowTuning to other tasks such as legal domain QA Zhong et al. (2020); Lyu et al. (2022, 2023a) and mathematical reasoning Luo et al. (2023). Moreover, our efforts have been concentrated on enhancing the knowledge awareness of LLMs during the fine-tuning stage. Future studies will aim to explore improving knowledge awareness of LLMs in the pre-training stage Rosset et al. (2020).
Ethical Considerations
KnowTuning mainly focuses on completeness, factuality, and logicality, but not social bias Pitoura et al. (2017); Lyu et al. (2023b) or the potential for generating harmful or toxic content Song et al. (2024); Hewitt et al. (2024); Gao et al. (2024). We plan to adopt our method to reduce social bias and harmful content at fine-grained and coarse-grained levels in future work.
Acknowledgments
This work was supported by the Natural Science Foundation of China (62272274, 62372275, 62102234, 62202271, 62072279), the National Key R&D Program of China with grant No.2022YFC3303004, the Natural Science Foundation of Shandong Province (ZR2021QF129), the China Scholarship Council under grant number 202306220180, the Dutch Research Council (NWO), under project numbers 024.004.022, NWA.1389.20.183, and KICH3.LTP.20.006, and the European Union’s Horizon Europe program under grant agreement No 101070212. All content represents the opinion of the authors, which is not necessarily shared or endorsed by their respective employers and/or sponsors.
@@ Appendix A Details of Datasets
  • NQ Kwiatkowski et al. (2019): We filter 200 questions and corresponding long answers as testing QA pairs from the development set. The length of these long answers ranges from 100 to 500.
  • ELI5 Fan et al. (2019): We filter 200 questions in the test set and the corresponding highest scoring answers as testing QA pairs.
@@ Appendix B Details of Baselines
  • FactTune Tian et al. (2023): We follow Min et al. (2023) to first break each candidate answers into individual facts, and prompt LLMs to measure the correctness of each fact based on the golden answer as a reference. [^2] https://github.com/shmsw25/FActScore Then, we construct factuality comparison sets by the percentage of correct facts. Finally, we adopt DPO Rafailov et al. (2023) for factuality comparison sets optimization.
@@ C.1 GPT-4 Evaluation
This section provides specifics of the GPT-4 prompt utilized for reference-based evaluation, employing gpt4-turbo. Figure 3 illustrates the adapted prompt from  Zheng et al. (2024a), aimed at assessing the completeness, factuality, and logicality of answers. To avoid positional bias Ko et al. (2020); Wang et al. (2023e), we evaluate each answer in both positions during two separate runs.
@@ C.3 Fine-grained facts evaluation
Following Min et al. (2023), we first break candidate answers into individual facts, and use gpt-3.5-turbo to measure the correctness of each fact based on the golden answer as a reference.2
@@ D.3 Training
During the training phase, the AdamW optimizer Loshchilov and Hutter (2019) is utilized with initial learning rates of $5\cdot 10^{-5}$ for SFT and $1\cdot 10^{-5}$ for DPO. The batch sizes for SFT and DPO are set to 32 and 16, respectively, with SFT undergoing 3 epochs of training and DPO 1 epoch. The filtering and deleting percentages, $\alpha$ and $\beta$, are both fixed at 0.5. The scalar weighting hyperparameter $\gamma$ is set to 0.2. We determine the hyperparameters through pilot experiments. Training leverages PEFT Mangrulkar et al. (2022), LLaMA-Factory Zheng et al. (2024b) and LoRA Hu et al. (2022).